UT 2012-07-29
=============

CHARA/CLIMB

B1   B2   B3   B4   B5   B6
S2   E2   W1
     E1   W2

S1: POP 5
S2: POP 5
E1: POP 1
E2: POP 1
W1: POP 4
W2: POP 5

Seeing: poor (r_0=5 cm)


*** S2-E2-W1

UT 04:29 - hd144585 (CAL for AS205A)
   	   recording LONG, 750 Hz, DEST

UT 05:20 - AS205A
   	   doing NIRO alignment

   	   recording LONG, 500 Hz, DEST
	   all three fringes clearly there, but they jump a lot (high airmass, 30 deg)
	   good data
   	   recording LONG, 500 Hz, NON-DES

UT 05:58 - hd145917 (CAL for AS205A)
   	   faint!  can only find one fringe, which is noisy

UT 06:15 - hd144585 (CAL for AS205A)
   	   saturates in NON-DEST, put in aperture wheel 1/4
   	   recording LONG, 500 Hz, NON-DES
	   take out aperture wheel
   	   recording LONG, 500 Hz, DEST (only Beam 1-2 = S2-E2 fringe there, elevation 25 deg)

UT 06:35 - hd158646 (CAL at similar elevation as AS205A)
	   (CHECK NAME AGAIN IN HEADER, LIKELY WRONG HD NUMBER)
   	   recording LONG, 500 Hz, DEST (only Beam 1-2 = S2-E2 fringe there, elevation 25 deg)
	   all fringes there, but beam 2-3 and 3-1 in the first 100 scans or so very bad; improves then

UT 06:55 - hd173113 (CAL for VV Ser)
   	   doing NIRO alignment
   	   recording LONG, 500 Hz, NONDEST, aperture wheel 1/2

UT 07:20 -VV Ser
   	   recording LONG, 500 Hz, NONDEST  (during the initial beam sequence, beam 1=S2 was lost)
	   all 3 fringes there, but low SNR
	   S2 tip-tilt drops out frequently

UT 08:00 - hd173113 (CAL for VV Ser)
   	   recording LONG, 500 Hz, NONDEST, aperture wheel 1/4
   	   
UT 08:20 - VV Ser
   	   doing NIRO alignment
   	   recording LONG, 500 Hz, NONDEST, abort since aperture wheel was still in
	   aperture wheel out
	   find only fringe on beam 2-3
   	   recording LONG, 500 Hz, NONDEST, lost beam 2-3 fringe, abort


*** S2-E1-W2 (POP 1-1-5)
change POP on S2 to 1

UT 09:00 - hd194708 (CAL for MWC340)
	   cannot find fringes

UT 09:30 - hd194193 (bright fringe finder)
   	   cannot find fringes
	   rehoming carts
	   W2 was off

UT 10:00 - another fringe finding star

UT 10:20 - hd199373 (CAL for V1057 Cyg)
   	   recording LONG, 500 Hz, NONDEST

UT 10:45 - V1057 Cyg
   	   problem with E2 during acquisition, have to give since won't make it before we run out of delay

UT 11:00 - hd216425
   	   recording LONG, 750 Hz, DEST

UT 11:15 - MWC1080
   	   problem with E1 acquisition system (cannot change contrast)
   	   find only the crossfringe
   	   recording LONG, 500 Hz, NONDEST, only beam 1-3 and noisy, maybe useless

UT 12:00 - closing down